Answer the following questions based on the demand curve: Q = 100 – 2P Determine the...

  1. Answer the following questions based on the demand curve: Q = 100 – 2P
  1. Determine the maximum value of Total Revenue
  2. Find the elasticity when revenue is maximized. Explain if you expected the answer.
  3. Find the equation of total revenue in terms of ‘Q’

Will a monopolist ever produce at a price less than 25? Explain.
Answer #1

Q = 100- 2P TR = PO = P(100-2P) TR = loop - 2P² TB = 100 - up dP put dTR dP = 0 100 - 4P = 0 100= up P = 25 - = - 4 so so d82- (-2)/25) = -1 50 1-1) = c) a = 100-2P 2P= 100-o P = 50 a TR = Po = (50-00 TR = 500 - 82 A monopolist always producer in the